# Marvel: Accelerating Safe Online Reinforcement Learning with Finetuned Offline Policy

### v2

@@ Appendix B Related Work
Online Safe RL.  Online safe RL approaches can be generally divided into several categories. The first category includes primal-dual based methods, such as PDO [21], which combines PPO [38] with the Lagrange multiplier method to obtain a policy that satisfies safety constraints. CPPO-PID [20] combines PID control with Lagrangian methods to dampen cost oscillations. Similar Lagrangian-based methods are applied in conjunction with other unconstrained safe RL algorithms, such as TRPO-lag, PPO-lag, and SAC-lag. CPO [7] inherits from TRPO [39], optimizing with the Lagrange multiplier method within the trust region. CUP [40] extends CPO by incorporating the generalized advantage estimator. In comparison, RCPO [41] uses different update rates for the primal and dual variables. Two-stage iterative methods have also been developed for online safe RL, e.g., PCPO [42] and FOCOPS [43]. Besides the primal-dual based methods, primal methods, which are also known as Lyapunov methods, have been leveraged in some studies for online safe RL. For instance, IPO [44] uses logarithmic barrier functions. P3O [45] employs an exact penalty function to derive an equivalent unconstrained objective and restrict policy updates within the trust region. [46] leverages Lyapunov functions to handle constraints, which contains two parts, safe policy iteration and safe value iteration. Additionally, some studies [47, 48] borrow techniques from the control theory, such as HJ reachability [49, 50] and control barrier functions [51], to ensure state-wise zero costs.
@@
Offline Safe RL.  Offline safe RL seeks to learn a safe policy from static datasets without online environmental interactions. Similar to online safe RL, Lagrangian methods can still be applied here, by adapting offline unconstrained RL algorithms like BCQ [52] and BEAR [37] to the safe RL setting. CPQ [19] uses a VAE to detect OOD [53] actions and penalizes them in terms of cost. COptiDICE [54] extends OptiDICE [55] by adding safety constraints and derives a safe policy through the stationary distribution of the optimal policy. FISOR [8] decouples the process of satisfying safety constraints from maximizing rewards and employs a diffusion model as the policy. VOCE [18] estimates Q-values of both cost and reward in a pessimistic way, mitigating extrapolation errors caused by OOD actions. Decision transformer (DT) [56] has also been applied to safe RL, leading to constrained decision transformer [57].
@@
O2O Unconstrained RL.  O2O RL has recently attracted much attention in the unconstrained case, where a policy pretrained on an offline dataset is used to assist online policy learning, e.g., through finetuning or serving as a guide policy. More specifically, [58, 59, 60] and [61] explore various combinations of offline demonstration data with online learning. The core idea is that pure offline RL often struggles with limited performance due to heavy reliance on dataset quality. However, if interaction with the environment is allowed, pffline pretrained policy can be finetuned for improved performance. However, naive implementation of this process often leads to suboptimal performance [12, 27]. AWAC [12] prioritizes actions with high advantage estimates, while AW-Opt [62] builds on AWAC by applying positive sample filtering and using hybrid actor-critic exploration during online finetuning. [63] finetunes the pretrained policy by balancing the offline and online datasets. FamO2O [13] trains a family of policies using a universal model and then employs a balance model to select the most suitable policy for each state. Cal-QL [64] constrains the updates to the Q-network during online finetuning to prevent underestimation of the Q-values. SO2 [35] improves Q-value estimation by updating Q-values more frequently and using noise-augmented actions. Instead of directly finetuning the pretrained policy, Jump-start RL [27] and PEX [36] follows another direction to leverage the offline policy, by using it to guide the update of the online policy during online learning.
@@ Appendix C Details on Baselines
Considering the characteristics of safe RL, which requires keeping the cost below a certain threshold, not all O2O unconstrained RL algorithms are suitable for O2O safe RL. For instance, AWAC [12], which maximizes the advantage function, has not yet been applied in the safe RL context. We compare Marvel with the following baselines:
SO2 [35]. By analyzing Q-value estimation in offline to online transitions, the SO2 algorithm achieves more accurate Q-value estimation through Perturbed Value Update and by increasing the frequency of Q-value updates.
JSRL [27]. JSRL employs an offline pretrained policy as the exploration policy and a policy under training during the online phase as the target policy. Initially, the exploration policy is used, followed by the target policy during online interaction to facilitate curriculum learning. To adapt to the safe RL setting, we update the Lagrange multipliers using the aPID method when updating the target policy.
PEX [36]. Similar to JSRL, PEX uses an offline pretrained policy and a policy under training during the online phase for online interaction. However, PEX selects one of the actions based on the Q-networks’s value estimation of actions chosen by the two policies. To meet the safe RL requirements concerning cost, like the modifications to JSRL, we use the aPID method to update the Lagrange multipliers.
@@
We selected SO2, JSRL, and PEX as baselines because they represent prominent methods in O2O RL, and adapting them to the safe RL context provides a meaningful comparison. Including these baselines allows us to demonstrate the effectiveness of Marvel in a fair and relevant context.
@@ D.1 More experiments
Additionally, we present the performance of the baseline algorithms and Marvel, along with the offline pretrained policy as the starting point, as shown in Fig. 3 and Fig. 6, and summarized in Table 2.
Table 2: In the table, the content following ”Offline” represents the performance of the pretrained policy, while the rest shows the results of online finetuning based on the pretrained policy using baseline methods and the Marvel algorithm.
(a) Result of Fig. 3
Environment Algorithm Reward Cost
BallCircle Offline 166.00 11.00
From Scratch 241.70 10.94
Warm Start 176.63 18.53
SO2 58.41 1.95
JSRL 1.62 165.39
PEX -4.27 39.85
Marvel (ours) 603.94 19.75
BallRun Offline 262.00 3.00
From Scratch 315.21 5.58
Warm Start 132.16 22.99
SO2 286.79 12.13
JSRL 174.04 92.10
PEX -555.58 88.84
Marvel (ours) 306.55 5.48
CarCircle Offline 265.00 14.00
From Scratch 115.20 12.73
Warm Start 141.55 24.05
SO2 115.91 9.07
JSRL 1.06 130.01
PEX -12.64 142.94
Marvel (ours) 341.32 19.45
CarRun Offline 544.00 72.00
From Scratch 293.50 7.20
Warm Start 391.40 16.90
SO2 522.23 16.22
JSRL 126.59 128.21
PEX 152.99 38.79
Marvel (ours) 547.52 18.20
AntCircle Offline 4.00 39.00
From Scratch 1.48 0.01
Warm Start 1.28 0.00
SO2 3.69 1.31
JSRL 0.56 1.91
PEX 6.93 2.27
Marvel (ours) 5.56 0.87
HalfCheetah Offline 113.00 17.00
From Scratch 1074.27 28.81
Warm Start 974.10 15.44
SO2 559.40 23.89
JSRL 27.55 3.83
PEX -155.56 0.18
Marvel (ours) 1543.17 20.31
(b) Result of Fig. 6
Environment Algorithm Reward Cost
DroneCircle Offline 32.00 6.00
From Scratch 0.77 28.81
Warm Start 0.63 18.25
SO2 136.84 17.49
JSRL 0.21 5.87
PEX 3.22 74.86
Marvel (ours) 117.54 22.19
AntRun Offline 42.00 0.00
From Scratch 30.87 0.11
Warm Start 12.16 0.18
SO2 59.40 0.00
JSRL 2.47 0.00
PEX 175.63 18.27
Marvel (ours) 61.05 0.00
Hopper Offline 62.00 15.00
From Scratch 467.07 35.19
Warm Start 234.03 8.44
SO2 730.94 5.28
JSRL 3.54 2.93
PEX 215.63 12.32
Marvel (ours) 310.82 7.09
Swimmer Offline 4.00 2.00
From Scratch 18.37 19.28
Warm Start 17.04 18.80
SO2 18.73 19.59
JSRL 0.11 0.00
PEX -12.98 0.00
Marvel (ours) 37.90 18.21
@@ D.2 More ablations
Figure 8: In the figure, “high” and “low” represent the different performance levels of offline pretrained policies resulting from varying quality in the offline dataset. These policies are then finetuned online. The results demonstrate that the Marvel algorithm is robust to both different offline dataset qualities and pretrained policy performances.
@@
D.3 Finetune Q-networks vs train new Q-networks in VPA
In Marvel, VPA fine-tunes the offline pretrained Q-networks. Fig. 9 illustrates the training curves when, instead of fine-tuning the pretrained Q-networks, the Q-networks are retrained from scratch during the VPA phase and subsequently fine-tuned online. As shown, fine-tuning the pretrained Q-networks achieves better performance. This is because, although the pretrained Q functions may be inaccurate, they still provide meaningful prior knowledge from the offline dataset and serve as a valuable starting point for Q function fine-tuning. This would generally speed up the learning and lead to a better local optima compared to learning from scratch based on the offline data from a random initial point. Moreover, considering the limited number of steps allowed in VPA for efficiency, directly learning completely forgoes the knowledge learned offline and can fail to find good Q estimations.
Figure 9: In the figure, “VPA (finetune)” refers to fine-tuning the offline pretrained Q-networks during the VPA phase, while “VPA (from scratch)” refers to training new Q-networks from scratch during the VPA phase.
D.4 Parameter sensitivity of aPID
D.4.1 PID
The SAC-lag algorithm in [34] utilizes PID control, with PID parameters carefully optimized. However, if their provided parameters are used directly under the environmental settings, policy updates, and Q-network update configurations of this paper, the performance is suboptimal. Fig. 10 presents a comparison, showing that when the PID parameters from the FSRL library are applied, the performance of online fine-tuning is significantly degraded. It is clear that the implementation of PID in our paper indeed significantly outperforms the implementation of PID provided by FSRL. More importantly, even with inappropriate PID parameters, aPID effectively boosts performance, achieving higher rewards while maintaining more stable cost levels.
Figure 10: “Our PID” and “Our aPID” refer to using the PID and aPID parameters proposed in this paper for adjusting the Lagrange multipliers, respectively. Similarly, “FSRL PID” and “FSRL aPID” represent the parameters provided by the FSRL library for the same purpose.
D.4.2 Parameters in aPID
@@
Figure 11: The three numbers within “[ ]” represent the three parameters for adaptively adjusting $K_{p}$, $K_{i}$ and $K_{d}$, namely $\alpha$, $\beta$ and $\gamma$.
D.5 Parameter sensitivity of VPA
@@
Figure 12: The two numbers inside “[ ]” represent the values of $\alpha$ and $\alpha_{c}$ used in VPA, as described in Eqs. 8 and 9.
D.6 Correctness of our implementation of SAC-lag
The primary goal of O2O safe RL algorithms is to achieve competitive performance with minimal environment interactions and in the shortest time by leveraging offline information to accelerate online learning. In contrast, the algorithm in [34] (and other similar online algorithms) achieves higher performance but relies on significantly more interactions. For example, in the BallCircle environment, [34] utilized 1.5 million environment interactions, whereas our method required only 120,000 interactions (with an average of 600 interactions per gradient update). This significant reduction highlights the efficiency of our approach, particularly in resource-constrained and safety-critical settings where the number of online interactions is strictly limited.
@@ D.6 Correctness of our implementation of SAC-lag
Figure 13: “Our SAC-lag” refers to the SAC-lag algorithm implemented in this paper, while “FSRL SAC-lag” represents the SAC-lag algorithm provided by the FSRL library. Using the same environment settings (including interaction steps) and update frequencies as in this paper, the results from the FSRL library are shown to be similar to ours.
D.7 Without VPA but with good initial values of the Lagrangian multipliers
As shown in Fig. 14, when VPA is not used and the Lagrange multipliers are updated using dual ascent (as described in Eq. 3), even with appropriately chosen initial values for the Lagrange multipliers (“Warm Start w/ lag init”), the performance, while better than initializing with zero (“Warm Start”), still falls short of achieving optimal results.
Figure 14: In the figure, “Marvel” and “Warm Start” follow the legend defined in Fig. 1. “Warm Start w/ lag init” represents the approach of empirically selecting appropriate initial values for the Lagrange multipliers and performing online fine-tuning.
D.8 Performance of O2O RL without considering safe RL
In the O2O safe RL setting, when the O2O unconstrained RL algorithm is applied, the online fine-tuning results are shown in Figure Fig. 15. As can be seen, the unconstrained RL algorithm, which focuses solely on maximizing reward without controlling the cost, achieves a high reward but results in a cost that exceeds the threshold, thereby violating the constraint. This experiment further demonstrates the necessity of incorporating safe settings in such environments.
Figure 15: In the figure, the labels are consistent with those in Fig. 3.

### v3

@@ Appendix B Related Work
Online Safe RL.  Online safe RL approaches can be generally divided into several categories. The first category includes primal-dual based methods, such as PDO (Chow et al., 2018a), which combines PPO (Schulman et al., 2017) with the Lagrange multiplier method to obtain a policy that satisfies safety constraints. CPPO-PID (Stooke et al., 2020) combines PID control with Lagrangian methods to dampen cost oscillations. Similar Lagrangian-based methods are applied in conjunction with other unconstrained safe RL algorithms, such as TRPO-lag, PPO-lag, and SAC-lag. CPO (Achiam et al., 2017) inherits from TRPO (Schulman, 2015), optimizing with the Lagrange multiplier method within the trust region. CUP (Yang et al., 2022) extends CPO by incorporating the generalized advantage estimator. In comparison, RCPO (Tessler et al., 2018) uses different update rates for the primal and dual variables. Two-stage iterative methods have also been developed for online safe RL, e.g., PCPO (Yang et al., 2020) and FOCOPS (Zhang et al., 2020). Besides the primal-dual based methods, primal methods, which are also known as Lyapunov methods, have been leveraged in some studies for online safe RL. For instance, IPO (Liu et al., 2020) uses logarithmic barrier functions. P3O (Zhang et al., 2022) employs an exact penalty function to derive an equivalent unconstrained objective and restrict policy updates within the trust region. (Chow et al., 2018b) leverages Lyapunov functions to handle constraints, which contains two parts, safe policy iteration and safe value iteration. Additionally, some studies (Wabersich et al., 2023; Choi et al., 2020) borrow techniques from the control theory, such as HJ reachability (Bansal et al., 2017; Yu et al., 2022) and control barrier functions (Ames et al., 2019), to ensure state-wise zero costs.
@@
Offline Safe RL.  Offline safe RL seeks to learn a safe policy from static datasets without online environmental interactions. Similar to online safe RL, Lagrangian methods can still be applied here, by adapting offline unconstrained RL algorithms like BCQ (Fujimoto et al., 2019) and BEAR (Kumar et al., 2019) to the safe RL setting. CPQ (Xu et al., 2022) uses a VAE to detect OOD (Ren et al., 2019) actions and penalizes them in terms of cost. COptiDICE (Lee et al., 2022a) extends OptiDICE (Lee et al., 2021) by adding safety constraints and derives a safe policy through the stationary distribution of the optimal policy. FISOR (Zheng et al., 2024) decouples the process of satisfying safety constraints from maximizing rewards and employs a diffusion model as the policy. VOCE (Guan et al., 2024) estimates Q-values of both cost and reward in a pessimistic way, mitigating extrapolation errors caused by OOD actions. Decision transformer (DT) (Chen et al., 2021) has also been applied to safe RL, leading to constrained decision transformer (Liu et al., 2023c).
@@
O2O Unconstrained RL.  O2O RL has recently attracted much attention in the unconstrained case, where a policy pretrained on an offline dataset is used to assist online policy learning, e.g., through finetuning or serving as a guide policy. More specifically, (Hester et al., 2018; Nair et al., 2018; Rajeswaran et al., 2017) and (Rudner et al., 2021) explore various combinations of offline demonstration data with online learning. The core idea is that pure offline RL often struggles with limited performance due to heavy reliance on dataset quality. However, if interaction with the environment is allowed, pffline pretrained policy can be finetuned for improved performance. However, naive implementation of this process often leads to suboptimal performance (Nair et al., 2020; Uchendu et al., 2023). AWAC (Nair et al., 2020) prioritizes actions with high advantage estimates, while AW-Opt (Lu et al., 2022) builds on AWAC by applying positive sample filtering and using hybrid actor-critic exploration during online finetuning. (Lee et al., 2022b) finetunes the pretrained policy by balancing the offline and online datasets. FamO2O (Wang et al., 2024) trains a family of policies using a universal model and then employs a balance model to select the most suitable policy for each state. Cal-QL (Nakamoto et al., 2024) constrains the updates to the Q-network during online finetuning to prevent underestimation of the Q-values. SO2 (Zhang et al., 2024) improves Q-value estimation by updating Q-values more frequently and using noise-augmented actions. Instead of directly finetuning the pretrained policy, Jump-start RL (Uchendu et al., 2023) and PEX (Zhang et al., 2023a) follows another direction to leverage the offline policy, by using it to guide the update of the online policy during online learning. Zhou et al. (2024) propose to finetune an offline policy without retaining the offline data. Their method first warms up training by interacting with the environment using the offline policy before performing any policy updates.
@@ Appendix C Details on Baselines
Considering the characteristics of safe RL, which requires keeping the cost below a certain threshold, not all O2O unconstrained RL algorithms are suitable for O2O safe RL. For instance, AWAC (Nair et al., 2020), which maximizes the advantage function, has not yet been applied in the safe RL context. We compare Marvel with the following baselines:
SO2 (Zhang et al., 2024). By analyzing Q-value estimation in offline to online transitions, the SO2 algorithm achieves more accurate Q-value estimation through Perturbed Value Update and by increasing the frequency of Q-value updates.
JSRL (Uchendu et al., 2023). JSRL employs an offline pretrained policy as the exploration policy and a policy under training during the online phase as the target policy. Initially, the exploration policy is used, followed by the target policy during online interaction to facilitate curriculum learning. To adapt to the safe RL setting, we update the Lagrange multipliers using the aPID method when updating the target policy.
PEX (Zhang et al., 2023a). Similar to JSRL, PEX uses an offline pretrained policy and a policy under training during the online phase for online interaction. However, PEX selects one of the actions based on the Q-networks’s value estimation of actions chosen by the two policies. To meet the safe RL requirements concerning cost, like the modifications to JSRL, we use the aPID method to update the Lagrange multipliers.
Cal-QL (Nakamoto et al., 2023). This method fine-tunes policies using a better-calibrated Q-function. We employ the aPID method to update the Lagrange multipliers.
SAC with an offline replay buffer. This baseline naively incorporates offline data into the online replay buffer. As with our modifications before, we use the aPID method to update the Lagrange multipliers.
@@
We selected SO2, JSRL, PEX and Warm Start as baselines because they represent prominent methods in O2O RL, and adapting them to the safe RL context provides a meaningful comparison. Including these baselines allows us to demonstrate the effectiveness of Marvel in a fair and relevant context.
@@ D.1 More experiments
In Fig. 4, we provide training curve of Marvel and baseline algorithms in BallRun, BallCircle, CarRun, CarCircle, HalfCheetah, AntCircle. In environments like BallCircle and CarCircle, Marvel finds a good policy within less than $15$ steps, dominating baseline methods in both performance and speed. Note that all approaches indeed start from the offline policy and Q-functions, i.e., the same point at step 0 (ignored in all figures).
@@
In Fig. 6, Cal-QL, which fine-tunes policies using a better-calibrated Q-function, demonstrates reasonable performance. However, since it was not originally designed for safe reinforcement learning, its direct application in this setting is inappropriate, since it fails to fully address the issues of erroneous Q-estimations and Lagrangian mismatch, as illustrated before. Similarly, the naive use of offline data, SAC with an offline replay buffer, fails to achieve satisfactory outcomes, as its incurred cost substantially exceeds the predefined threshold. This violation of safety constraints highlights the algorithm’s inability to succeed under such conditions.
@@
Figure 6: This figure presents a performance comparison between Marvel, Cal-QL, and SAC with an offline replay buffer. As shown, Marvel achieves the best overall performance.
In Table 5, We conpare our Marvel with CDT(Liu et al., 2023c). Due to CDT’s design for offline settings, we use the replay buffer to apply CDT in the online setting, which is represented as "CDT-finetune" in the table. It can be seen that CDT, having more parameters, generally performs better than CPQ in the offline phase. However, since the finetuning process was not specifically designed, performance stagnation and even degradation occur during finetuning.
Table 5: Comparison of Marvel and CDT
Environment Offline Online
CPQ CDT Marvel (ours) CDT-finetune
BallCircle 166.0 / 11.0 521.7 / 18.3 603.9 / 19.8 468.7 / 9.3
BallRun 262.0 / 3.0 448.5 / 48.5 306.6 / 5.5 449.0 / 50.3
CarCircle 265.0 / 14.0 328.7 / 18.1 341.3 / 19.5 320.2 / 23.1
CarRun 544.0 / 72.0 553.9 / 25.1 547.5 / 18.2 542.0 / 23.3
AntCircle 4.0 / 39.0 221.6 / 54.7 5.6 / 0.9 161.2 / 36.4
HalfCheetah 113.0 / 17.0 1358.0 / 24.3 1544.8 / 20.0 1536.7 / 3.9
@@ D.2.1 VPA on Q vs Qc vs Both and if VPA need entropy term
D.2.2 Finetune Q-networks vs train new Q-networks in VPA
In Marvel, VPA finetunes the offline pretrained Q-networks. Fig. 8 illustrates the training curves when, instead of finetuning the pretrained Q-networks, the Q-networks are retrained from scratch during the VPA phase and subsequently finetuned online. As shown, finetuning the pretrained Q-networks achieves better performance. This is because, although the pretrained Q functions may be inaccurate, they still provide meaningful prior knowledge from the offline dataset and serve as a valuable starting point for Q function finetuning. This would generally speed up the learning and lead to a better local optima compared to learning from scratch based on the offline data from a random initial point. Moreover, considering the limited number of steps allowed in VPA for efficiency, directly learning completely forgoes the knowledge learned offline and can fail to find good Q estimations.
Figure 8: In the figure, “VPA (finetune)" refers to finetuning the offline pretrained Q-networks during the VPA phase, while “VPA (from scratch)" refers to training new Q-networks from scratch during the VPA phase.
D.2.3 Marvel’s Ability to Improve Pretrained Policies at Different Performance Levels
@@
Figure 9: In the figure, “high" and “low" represent the different performance levels of offline pretrained policies resulting from varying quality in the offline dataset. These policies are then finetuned online. The results demonstrate that the Marvel algorithm is robust to both different offline dataset qualities and pretrained policy performances.
D.3 Parameter sensitivity of aPID
D.3.1 PID
The SAC-lag algorithm in (Liu et al., 2024) utilizes PID control, with PID parameters carefully optimized. However, if their provided parameters are used directly under the environmental settings, policy updates, and Q-network update configurations of this paper, the performance is suboptimal. Fig. 10 presents a comparison, showing that when the PID parameters from the FSRL library are applied, the performance of online finetuning is significantly degraded. It is clear that the implementation of PID in our paper indeed significantly outperforms the implementation of PID provided by FSRL. More importantly, even with inappropriate PID parameters, aPID effectively boosts performance, achieving higher rewards while maintaining more stable cost levels.
Figure 10: “Our PID" and “Our aPID" refer to using the PID and aPID parameters proposed in this paper for adjusting the Lagrange multipliers, respectively. Similarly, “FSRL PID" and “FSRL aPID" represent the parameters provided by the FSRL library for the same purpose.
D.3.2 Parameters in aPID
@@
Figure 11: The three numbers within “[ ]" represent the three parameters for adaptively adjusting $K_{p}$, $K_{i}$ and $K_{d}$, namely $\alpha$, $\beta$ and $\gamma$.
D.4 Parameter sensitivity of VPA
@@
Figure 12: The two numbers inside “[ ]" represent the values of $\alpha$ and $\alpha_{c}$ used in VPA, as described in Eq. 8 and Eq. 9.
D.5 Correctness of our implementation of SAC-lag
The primary goal of O2O safe RL algorithms is to achieve competitive performance with minimal environment interactions and in the shortest time by leveraging offline information to accelerate online learning. In contrast, the algorithm in (Liu et al., 2024) (and other similar online algorithms) achieves higher performance but relies on significantly more interactions. For example, in the BallCircle environment, (Liu et al., 2024) utilized 1.5 million environment interactions, whereas our method required only 120,000 interactions (with an average of 600 interactions per gradient update). This significant reduction highlights the efficiency of our approach, particularly in resource-constrained and safety-critical settings where the number of online interactions is strictly limited.
@@ D.5 Correctness of our implementation of SAC-lag
Figure 13: “Our SAC-lag" refers to the SAC-lag algorithm implemented in this paper, while “FSRL SAC-lag" represents the SAC-lag algorithm provided by the FSRL library. Using the same environment settings (including interaction steps) and update frequencies as in this paper, the results from the FSRL library are shown to be similar to ours.
D.6 Without VPA but with good initial values of the Lagrangian multipliers
As shown in Fig. 14, when VPA is not used and the Lagrange multipliers are updated using dual ascent (as described in Eq. 3), even with appropriately chosen initial values for the Lagrange multipliers (“Warm Start w/ lag init"), the performance, while better than initializing with zero (“Warm Start"), still falls short of achieving optimal results.
Figure 14: In the figure, “Marvel" and “Warm Start" follow the legend defined in Fig. 1. “Warm Start w/ lag init" represents the approach of empirically selecting appropriate initial values for the Lagrange multipliers and performing online finetuning.
D.7 Additional Baseline: Distilling Q-values of the Behavior Policy
To further evaluate the effectiveness of Marvel, we include an additional baseline that distills the Q-values of the policy which generated the offline dataset. In offline RL, the dataset is not necessarily produced by a single policy; it can be a mixture of trajectories collected from multiple behavior policies of varying quality. To align with such a mixed dataset, we first use behavior cloning (BC) to obtain a policy consistent with the offline data distribution. We then perform off-policy evaluation to distill the Q-values of this cloned policy and use the resulting critics to initialize online finetuning.
However, our experiments show that this approach fails to keep the cost below the safety threshold during online finetuning. Because the BC policy directly imitates actions from the offline dataset, it inevitably learns unsafe action patterns present in the data. When the distilled critics are subsequently used for online updates, these unsafe patterns remain and cannot be corrected even with the adaptive control provided by aPID. As a result, the policy cannot maintain safety and exhibits performance significantly worse than MARVEL.
Figure 15: Comparison between Marvel and the additional baseline that distills Q-values of the behavior policy. The BC-distilled baseline fails to control the cost below the threshold due to inherited unsafe patterns from the offline dataset, leading to high constraint violations and lower returns, whereas MARVEL effectively maintains safety and achieves superior performance.